OpenStudy (anonymous):

Which of the following functions is/are even? (I) g (t) = t^4 − 45 (II) h (t) = t^4 + t^3 (III) k (t) = t^4 + t^2 + 120

OpenStudy (ksaimouli):

i think 111

OpenStudy (ksaimouli):

because all are even (^42 and 120)

OpenStudy (anonymous):

how do i determine whats even or not

OpenStudy (ksaimouli):

if the power is even then the whole thing is even

OpenStudy (hunus):

\[f(-x)=f(x)\] A function is even if it obeys this formula. What this means is that the graph is symmetrical about the y axis. For instance, cosine is an even function because \[f(-\pi)=f(\pi)\] \[-1 = -1\]